Rick earns $500 less than 3 times as much as Jim. If their combined salary is $49,500, how much do they each earn?
Answer by checkley77(12844)

LET (3X-500) BE RICK'S EARNINGS.
LET X BE JIM'S EARNINGS.
X+3X-500=49,500
4X=49,500+500
4X=50,000
X=50,000/4
X=12,500 JIM'S EARNINGS.
3*12,500-500
37,500-500=37,000 RICK'S EARNINGS.
PROOF:
12,500+37,000=49,500
49,500=49,500